### v1

# Extremely high-density ($>10^{14}$cm-2) with very low sheet resistance 2D electron gases in N-polar AlGaN/GaN heterostructures with GaN/AlN superlattice back barriers grown on sapphire substrates

Abstract
We report on the realization of extremely high-density ($>10^{14}$cm-2) 2D electron gases in N-polar AlGaN/GaN heterostructures grown on sapphire substrates. By introduction the GaN/AlN superlattice (SL) back barrier between GaN buffer layer and AlGaN barrier layer we observed a giant enhancement of the 2D electron gas density at the GaN/Al0.3Ga0.7N interface from $3\times 10^{13}$cm-2 (without SL) to $1.4\times 10^{14}$cm-2 (with SL back barrier) that is only an order of magnitude below the intrinsic crystal limit of $\approx 10^{15}$cm-2. We found that the changes of 2D electron gas density with SL correlated well with changes of the wafer warp parameter which suggested the strain-induced 2D electron gas density. Simultaneously, the room temperature electron mobility was as high as 169 cm2/Vs, which with the electron density of $1.4\times 10^{14}$cm-2 gives a very low sheet resistance of 264 $\Omega$/sq (one of the lowest reported so far for the N-polar 2D electron gas channel). The finding of the very high-density 2D electron gas with very low sheet resistance in N-polar III-nitride heterostructures can push the power performance N-polar GaN HEMTs into a new level which is unreachable by the conventional N-polar GaN HEMTs.
preprint: AIP/123-QED
I Introduction
The discovery of the high-conductivity quantum-confined two-dimensional (2D) electron gases at the interface of AlGaN/GaN semiconductor heterostructures in the mid-1990s[23] has generated considerable attention in the scientific community since its formation did not require the presence of chemical dopants. Contrary to AlGaAs/GaAs material systems[2, 3, 4], 2D electron gases in AlGaN/GaN heterostructures are formed spontaneously due to net polarization charges composed of spontaneous and piezoelectric polarization of AlGaN and GaN.[5, 6]
In the context of the high electron mobility transistors (HEMTs) application a high-2D electron gas density together with high electron mobility is strongly desired[7]. In particular, the high-2D electron density reduces the parasitic series resistances in the source/drain which leads to increasing of the on-currents, transconductance and in consequence power output. In addition, a high-2D electron density is needed to reduce the effect of surface and buffer traps during high frequency and high voltage operation.[8, 9] Therefore, by engineering the 2D electron density the performance of AlGaN/GaN HEMTs can be greatly improved. Beside HEMTs applications, very high-2D electron density systems are also attractive for infrared plasmonic applications[10, 11] or “charge gain” devices.[12, 13]
Recently, the N-polar III-nitride heterostructures attracted great attention in nitride community due to their several advantages over metal-polar counterpart in high frequency and high-power applications.[14, 15, 16, 17, 18, 19, 20] The typical as-grown 2D electron gas density in N-polar III-nitride heterostructures ranging from low-$6\times 10^{12}$ to high $2\times 10^{13}$cm-2 with mobility between 900–1500 cm2/Vs and sheet resistances from 300 to 650 $\Omega$/sq.[14, 21, 22, 23, 24, 25] However, it should be highlighted that the highest observed 2D electron gas densities in N-polar III-nitride heterostructures were $4.3\times 10^{13}$cm-2 with mobility 450 cm2/Vs and sheet resistance 320 $\Omega$/sq.[21, 22] Such gases were realized using a high Al-content AlGaN barrier and AlN substrates.[21, 22] The 2D electron densities higher than $10^{14}$cm-2 have not been measured in N-polar III-nitride semiconductor heterostructures so far in an as-grown structure or field-effect induced channel.
Figure 1: (a) Schematic illustration of the fabricated epitaxial structures used in this study. (b) Root mean square (RMS) roughness as a function of SL periods. Inset of (a) shows the TEM image of the AlGaN/GaN heterostructure with 4 period GaN/AlN SL structure. Inset of (b) shows the 2D and 1D surface profiles from interferometric surface profilers of the AlGaN/GaN heterostructure with 4 period GaN/AlN SL
In this work, the extremely high-2D electron densities ($>10^{14}$cm-2) are demonstrated in N-polar AlGaN/GaN heterostructure. It was found that introducing the GaN/AlN superlattice (SL) back barrier between GaN buffer layer and AlGaN barrier layer leads to a giant enhancement of 2D electron gas density at GaN/AlGaN interface. In particular, we observed an increase of 2D electron gas density from $3\times 10^{13}$cm-2 (without SL structure) to $1.4\times 10^{14}$cm-2 (with SL) that is only an order of magnitude below the intrinsic crystal limit of $\approx 10^{15}$cm-2. The Hall measurement shows that this high-2D electron gas density has a room temperature electron mobility as high as 169 cm2/Vs which with 2D electron gas density of $1.4\times 10^{14}$cm-2 leads to a very low sheet resistance of 264 $\Omega$/sq.
Figure 2: (a) Room-temperature 2D electron gas density as a function of the SL period number, (b) room-temperature electron mobility vs. SL period number and (c) room-temperature Hall sheet resistance as a function of the SL period number. Inset of (a) shows schematically illustration of the warp parameter.
Figure 3: Distribution of polarization charges in the investigated SL based N-polar AlGaN/GaN heterostructures.
II SAMPLE FABRICATION
Fig. 1(a) shows the schematic illustration of the fabricated N-polar AlGaN/GaN heterostructures. The investigated structures were grown by Metal-Organic Chemical Vapour Deposition (MOCVD) using trimethylgallium (TMGa), trimethylaluminum (TMAl), and ammonia as precursors on (0001) sapphire with a misorientation angle of 4o toward the a-sapphire-plane. Firstly, a 100 nm AlN layer was grown at 1025 oC with a rate of 2 $\mu$m/h. Subsequently, a 400 nm unintentionally doped GaN buffer layer was grown at 1025 oC with a V/III ratio of 1600 at the pressure of 20 kPa. On top of the buffer GaN layer a thin superlattice (SL) structure was deposited, which contains the alternating GaN and AlN layers (see inset of Fig. 1(a)). The thickness of GaN and AlN layers in SL structure was 10 and 5 nm, respectively. The number of SL periods, n, (inset of Fig. 1(a)) was varying between 0 to 10. Next, a 30-nm-thick Al0.3Ga0.7N barrier layer was deposited followed by a 20 nm thick undoped GaN channel layer. Finally, 3-nm-thick Al0.4Ga0.6N and 2-nm-thick GaN cap layers were grown on the top of the structure. The polarity of the fabricated N-polar AlGaN/GaN heterostructures was confirmed by the KOH etching and X-ray diffraction (XRD) measurements (not shown). The transmission electron microscope (TEM) image confirmed that the AlN/GaN SL structure was successfully grown (see inset of Fig. 1(a)). The surface morphologies of the fabricated N-polar AlGaN/GaN heterostructures were measured using an interferometric surface profiler. The measurement from the large area over 400 $\mu$m 300 $\mu$m indicates the smooth morphology with a root mean square ($RMS$) roughness between 4.2 to 4.5 nm for 0 to 6 period SL, as shown in Fig. 1(b). However, above 6 period SL, the surface becomes very rough with $RMS$ value of 8.5 nm for 10 period SL (Fig. 1(b)).
Figure 4: Simulated (a) band diagram and (b) electron density of AlGaN/GaN heterostructure with 4 period GaN/AlN SL structure. Simulated (c) band diagram and (d) electron density of the control sample, i.e., 4-period (GaN/AlN) SL/GaN (400 nm)/AlN (100 nm)/sapphire.
III RESULTS AND DISCUSSION
To characterize the electron transport properties in the fabricated N-polar AlGaN/GaN heterostructures, we applied the Hall-effect measurements with the Van der Pauw method using the aluminum-titanium Ohmic contacts and the magnetic field of 1 T. Fig. 2(a) and (b) show the measured room temperature 2D electron density and mobility as a function of the SL period number. Without SL structure the N-polar AlGaN/GaN heterostructures exhibits the 2D electron gas density of $3\times 10^{13}$cm-2 with mobility of 440 cm2/Vs. The mobility value is similar to the previously reported one for N-polar III-nitride heterostructures with the electron density of $3-4\times 10^{13}$cm-2.[21, 22] By increasing the SL period number, the 2D electron gas density initially slowly increases and then jumps rapidly at 4 period SL reaching the concentration of $1.4\times 10^{13}$cm-2 (see Fig. 2(a)). Further, the increasing of SL periods leads to the saturation of the electron gas density at the level $1.4\times 10^{13}$cm-2. Simultaneously, the electron mobility decreases with increasing of SL periods up to 8 and then becomes constant (Fig. 2(b)). We attributed this effect to the interface roughness scattering mechanism. In particular, Singisetti et al.[24] showed that the interface roughness is a dominant scattering mechanism limiting the electron mobility in the N-polar GaN quantum well channels. On the other hand, the interface roughness scattering is proportional to the square of the electric field which is likely to increase in the region of GaN/AlGaN interface with increasing the SL period number. Fig. 2(c) shows the room temperature Hall sheet resistance ($R_{H}$) as a function of the SL period number. As can be seen, $R_{H}$ exhibits a U-shaped dependency on the SL period number, namely with increasing n, $R_{H}$ gradually decreases reaching the minimum at 4 period SL and subsequently increases. In the range from 0 to 4 period SL, RH decreases almost 2 times which is a direct consequence of the rapid increasing of the 2D electron gas density (Fig. 2(a)). On the other hand, in the range from 4 to 8 period SL, RH increases approximately 4.5 times. This increases is due to the fact that the 2D electron density is practically saturated in the range from 4 to 8 period SL. However, the electron mobility in this range still continues decreasing from 169 at 4 period SL to 80 cm2/Vs at 8 period SL (Fig. 2(b)).
To explain the above results, we performed the Technology Computer Aided Design (TCAD) simulations of the band diagram of investigated heterostructures. In the simulation, we assumed the polarization charges ($\sigma$) at AlN/GaN, Al0.3Ga0.7N/GaN and Al0.4Ga0.6N/GaN interfaces given by the following equation:[5]
$\frac{\sigma}{q}=6.41\times 10^{13}-1.17\times 10^{13}x(1-x)$ (1)
where $\sigma$ is in the unit of cm-2 and $x$ is the Al composition. In addition, we assumed that on the GaN surface the positive spontaneous polarization charge of $\sigma_{GaN}$=$1.8\times 10^{13}$cm-2 is compensated by the surface negative charge $\sigma_{S}$=-$1.8\times 10^{13}$cm-2. Such compensation charge was previously found on the N-polar GaN surface by the X-ray photoelectron spectroscopy measurements[26]. The detailed distribution of the assumed polarization charges in the heterostructure is shown in Fig. 3.
The band diagram simulation clearly shows that in addition to the 2D electron channel at the GaN/AlGaN interface, the high-density 2D electron gases should be formed in the SL structure (see Figs. 4(a) and (b)). The formation of the parallel 2D electron channels in the SL structure could explain a significant decreasing of RH between 0 and 4 period SL (Fig. 2(a)). However, the RH increasing between 4 to 10 period SL is the signature that 2D electron gases are not formed in the SL structure (with an increase of the number of SL periods more 2D electron channels are formed in the SL structures and thus we should observe decreasing of the RH value). To confirm the lack of 2D electron gases in the SL structure we fabricated the control sample, i.e. 4 period SL/GaN (400 nm)/AlN (100 nm) /sapphire. The band diagram simulations indicated that the SL structure in the control sample should contain the similar very high density-2D electron gas as investigated ones (see Figs. 4(c) and (d)). However, the Hall-effect measurements from the control sample shows high contactless sheet resistance of 710 Ohm/s, which is similar to the sheet resistance of the 400 nm N-polar GaN layer (800 Ohm/sq). This means that 2D electron gas was not formed in the SL structure. One of the reasons why the 2D electron gas in the SL structure is not observed may be the fact that AlN in the SL is not pure AlN but a high Al-content AlGaN. For example, if instead of the AlN layer the Al0.6Ga0.4N layer was formed in SL the 2D electron gas should not appear, as shown in Figs. 5(a) and (b). In light of the lack of 2D electron gases in the SL structure, we can assume that the carrier transport in the investigated heterostructures (Fig. 1(a)) was mainly caused by the 2D electron gas at the GaN/Al0.3Ga0.7N interface. However, it should be noted that besides the 2D electron channel at the GaN/Al0.3Ga0.7N interface there is also some 2D electron channel in the GaN cap layer (see Figs. 5(a) and (b)). According to the TCAD simulation this 2D electron gas has a negligible concentration compared to the 2D electron gas at the GaN/Al0.3Ga0.7N interface (Fig. 5(b)). Thus, it can be concluded that the carrier transport was entirely dominated by the 2D electron gas at the GaN/Al0.3Ga0.7N interface.
Figure 5: Simulated (a) band diagram and (b) electron density of AlGaN/GaN heterostructure with 4 period GaN/Al0.6Ga0.4N SL structure.
Figure 6: (a) Schematic illustration of warp parameter reduction in the investigated epitaxial structures. (b) Directions of the spontaneous and piezoelectric polarizations in the examined structures. Enhancement of 2D electron density at the GaN/Al0.3Ga0.7N interface occurred due to reduction of $P_{PE}$ in GaN channel after SL deposition.
After excluding the formation of 2D electron gas in the SL structure another reason for enhancement of 2D electron gas density needs to be identify. The enhancement of the stress due to SL deposition is a natural factor that could cause increasing of the 2D electron gas. For example, previously K.-S. Im et al.[27] reported a giant formation of 2D electron gas density $>10^{14}$cm-2 in the metal-polar AlGaN/GaN heterostructure on Si substrates by generation of a strong tensile stress. To estimate roughly the strain in our heterostructures we measured the standardized warp parameter (flatness of the wafer) as shown in Fig. 2(a). 6. One can note that the dependencies of the warp parameter as a function of the SL period number is strikingly similar to the dependence of 2D electron gas vs. SL period number. In particular, the biggest changes of the warp parameter occur between 2 and 4 period SL, where the giant enhancement of the 2D electron gas density takes place (Fig. 2(a)). The reduction of the warp parameter in the range between 2 and 4 period SL indicates that the wafer becomes more flatten. On the other hand, the bow parameter was positive which means that the curvature of the wafer was concave. Thus the wafer flattening occurred probably due to the generation of compressive stress in the SL structure, which reduced the tensile strain, as schematically shown in Fig. 6(a). The reduction of the tensile strains can lead to enhancement of the 2D electron density at the GaN/Al0.3Ga0.7N interface in the following manner. Firstly, we recall the definition of the piezoelectric polarization $P_{PE}$ vector:
$P_{PE}=2\epsilon_{xx}(\frac{e_{31}-e_{33}C_{13}}{C_{33}})$ (2)
where $e_{31}$ and $e_{33}$ are the piezoelectric coefficients, $\epsilon_{xx}$ is the in-plane component of the strain tensor, and $C_{13}$ and $C_{33}$ are components of the compliance tensor.
Form above equation it follows that if the GaN layer is under the tensile strain, the $P_{PE}$ vector in the GaN channel has the same direction as spontaneous polarization ones $P_{SP}$ (see Fig. 6(b)) since the factor $\epsilon_{xx}(\frac{e_{31}-e_{33}C_{13}}{C_{33}})$$>$0. This means that $P_{PE}$ in the GaN layer according to Fig. 6(b) reduces the 2D electron gas density at the GaN/Al0.3Ga0.7N interface. On the other hand, as we mentioned before the tensile strain become reduced when SL is deposited (Fig. 6(a)). Thus, in the AlGaN/GaN heterostrucutres with SL, $P_{PE}$ in the GaN channel should be lower than in the structures without SL (see Fig. 6(b)). If we assumed now that decreasing of $P_{PE}$ in the Al0.3Ga0.7N barrier layer after deposition of SL is much lower in the case of the GaN channel (Fig. 6 (b)) then the enhancement of the 2D electron density at the GaN/Al0.3Ga0.7N interface will occur due to reduction of $P_{PE}$ in the GaN channel. However, it should be noted that there is a weak point of this explanation, namely the changes of $P_{PE}$ in the GaN channel itself may be too weak to induce such high 2D electron density (Fig. 2(a)). Thus, we believe that there are other stress-related contributions leading to the enhancement of the 2D electron density at GaN/Al0.3Ga0.7N interface. Another factor which could enhance the 2D electron density is that the introduced SL may act as excellent electron blocking layers suppressing the escape of electrons from the quantum well at the GaN/Al0.3Ga0.7N interface.
Figure 7: Room-temperature electron density and mobility reported in N polar III-nitride semiconductor heterostructures grown on various substrates: SiC, sapphire, GaN and AlN[14, 21, 22, 23, 24, 25].
Fig. 7 shows the benchmark plot comparing the 2D electron gas densities obtained in this work with the as-grown single-2D electron gases reported in the literature for N-polar III-nitride semiconductor heterostructures on various substrates: SiC, sapphire, GaN and AlN. The GaN/AlGaN 2D electron gases fabricated in this work exhibit the highest densities among all N-polar III-nitride heterostructures: they are more than three times larger than highest 2D electron gases observed so far in N-polar III-nitride heterostructures grown on AlN substrates and only an order of magnitude below the intrinsic crystal limit of $\approx 10^{15}$cm-2. Furthermore, in the case of 4 period SL, $R_{H}$ reach 269 Ohm/sq which is one of the lowest reported to date for N-polar 2D electron gas channel. The combination of such low RH and very high 2D electron gas density can lead to development of N-polar GaN HEMTs with even greater power performance.
IV CONCLUSIONS
In summary, we reported on the realization of the extremely high-density 2D electron gases in N-polar AlGaN/GaN heterostructures grown on sapphire substrates. We found that the introduction of the GaN/AlN superlattice back barrier between the GaN buffer layer and AlGaN barrier layer leads to a giant enhancement of 2D electron gas density at the GaN/Al0.3Ga0.7N interface from $3\times 10^{13}$cm-2 (without SL) to $1.4\times 10^{14}$cm-2. The enhancement of the 2D electron gas density was attributed to the increase of residual stress due to the SL deposition. Simultaneously, the room temperature electron mobility (at the 2D electron gas densities $1.4\times 10^{14}$cm-2) was as high as 169 cm2/Vs which led to the very low Hall sheet resistance of 264 Ohm/sq (one of the lowest reported so far for N-polar 2D electron gas channel). The realization of N-polar III-nitride heterostructures with very high 2D electron gas density and very low RH can push the power performance N-polar GaN HEMTs into a new level which the conventional N-polar GaN HEMTs is not able to reach.
Acknowledgements.
The authors express gratitude to Norikazu Nakamura for his kind support and discussions. The authors would like to thank Yoshiharu Kinoue and Takaaki Sakuyama for their support in experiments.
DATA AVAILABILITY
The data that support the findings of this study are available from the corresponding author upon reasonable request.